OpenStudy (anonymous):

Can some one please explain this to me in the simplest possible language? A straight fence is to be constructed from posts 6 inches wide and separated by lengths of chain 5 feet long. If a certain fence begins and ends with a post, which of the following could not be the length of the fence in feet? (12 inches = 1 foot) A. 17 B. 28 C. 35 D. 39 E. 50

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

to work out the total length :of 6 posts and 5 chains:= 6 * 0.5 = 3 ft this is the length of the 6 posts and we have 5 chains at 5 feet loing - total length %*5 = 25 ft This is a total length of 25 + 3 = 28 ft. So B could be a total length We can work out a formula for working out the length of a fence and that is L = 0.5x + 5y where x = number of posts and y = number of chains (0.5 ft and 5 ft) but we know that there must be 1 less chain than posts as there must be a post at each end. so y = x-1 And our formula becomes L = 0.5 x + 5(x-1) L = 5.5x - 5 - THats our formula Now look at the options 28 = 5.5 * 6 - 5 = 33 -5 = 28 so 28 ok could 17 be the length: 17 = 5.5x - 5 5.5x = 22 x = 4 Yes 17 is Ok could 35 be length 35 = 5.5x -5 5.5x = 40 x = 40/5.5 = 7.3 THIS IS THEONE WE ARE LOOKING FOR x (the number of posts) must of course must be a whole number. So C. 35 cannot be a length. I assume there is only one of these - you can try the other two to make sure if you like